For the next seven years, the St. Louis Cardinals will pay Albert Pujols an average of 14.7 million dollars per year. Twenty-five $20-bills weigh 0.9 ounces. If Albert were to insist on being paid in $20-bills, how many pounds would the $14.7 million weigh? Round to the nearest pound.
A. 1,654 B. 8,148 C. 26,460 D. 33,075 E. 41,344


Answer by Theo(13342)

25 * 20 = 500 dollars.

each set of 25 * 20 dollar bills is worth 500 dollars.

14.7 million divided by 500 = .0294 million sets of 25 * 20 dollar bills.

multiply that by 1 million to get 29400 sets of 25 * 20 dollar bills.

each set weighs .9 ounces.

total weight would be 29400 * .9 = 26460 ounces.

number of ounces in a pound is 16.

total weight would be 26460 / 16 = 1653.75 pounds.

round to the nearest pound to get 1654 pounds.

your answer would be selection A (1654 pounds).

algebraically, you would say:

let x = 1 set of 25 * 20 dollar bills.

500*x = value of each set.

.9*x = weight of each set.

your equation would be:

500*x = 14.7 million

divide both sides by 500 to get:

x = 14.7 milli9on / 500 = .0294 million * 1000000 = 29400

you would solve for the weight by saying

.9*x = .9*29400 = 26460 ounces.

since 16 ounces to a pound, you would then solve for pounds by saying.

26460/16 = 1653.75 pounds.

this rounds to 1654 pounds to the nearest pound.